Find the magnitude and direction (in degrees) of the vector.
v = 3i +√7j

Answer by jim_thompson5910(35256)

If v = ai+bj, then ||v|| = sqrt(a^2+b^2)


In this case, a = 3 and b = sqrt(7)


||v|| = sqrt(a^2+b^2)


||v|| = sqrt(3^2+(sqrt(7))^2)


||v|| = sqrt(9+7)


||v|| = sqrt(16)


||v|| = 4


So the magnitude of v is 4


To find the direction or angle, we use the idea that theta = arctan(b/a)


theta = arctan(b/a)


theta = arctan(sqrt(7)/3)


theta = 41.409622 (note: this is approximate)


So the direction is roughly 41.409622 degrees.